sir bansalquadratic eq page 24 q5  page 58 q 7 PROGRESSION PAGE #) THE2nd illustration seems wrong   binomial page 45 q15

Grade:11

1 Answers

Harsh Patodia IIT Roorkee
9 years ago
Pg 24 Q5
For 2nd Equation a2(b2-c2)x2 + b2(c2-a2)x + c2(a2-b2)=0
Clearly sum of coefficients is 0 i.e. 1 is the root

Since this has equal roots and one commot with 4x2sin2\theta – 4xsin\theta + 1
\Rightarrow 1 is root of this equation
\Rightarrow 4(1)2sin2\theta -4(1)sin\theta +1 =0
4 sin2\theta – 4sin\theta + 1=0
\Rightarrow (2sin\theta -1)2=0
or sin\theta =1/2
Cleary \theta = \pi/6 or 5\pi/6 satisfying the given condition.